A bomb of 18kg explodes into 12 kg and 6 kg the velocity of 12 kg is 5 m/s the KE of the other piece

Grade:12th pass

1 Answers

Mahima Kanawat
6 years ago
Dear student ....By momentum conservation , mv=m1v1+m2v2 , 18×0=12×5+6×v2 , on solving v2= 10 m/s. Therefore , KE =1/2mv²=1/2×6×100=300 J.....With Regards....Mahima....(AskIITians Forum Expert )